Total costs $80,944$141,626 Required: a. Calculate the missing costs. (Do not round intermediate calculations.) February August 8,400 S 24,360 Production level, in units Variable costs Fixed costs Mixed costs Total cost:s 20,160 30,700 25,884 S 80,944 $ 141,626 b. Calculate the cost formula for mixed cost using the high-low method. (Do not round intermediate calculations. Round Variable cost to 2 decimal places.) Cost formula per unit c. Calculate the total cost that would be incurred for the production of 12,676 units. (Do not round intermediate calculations.) Total costThe following data have been extracted from the records of Puzzle, Inc.: February August Production level, in units 8,400 20,160 Variable costs $ 24,360 $ ? Fixed costs ? 30,700 Mixed costs 25,884 ? Total costs $ 80,944 $ 141,626
Answer #1

Requirement a.

February

August

Production level in units

8400

              20,160

Variable cost

$   24,360.00

$    58,464.00

Fixed cost

$   30,700.00

$    30,700.00

Mixed cost

$   25,884.00

$    52,462.00

Total cost

$   80,944.00

$ 1,41,626.00

Working

February

August

Production level in units

8400

20160

Variable cost

24360

=2.9*20160

Fixed cost

30700

30700

Mixed cost

25884

=141626-58464-30700

Total cost

80944

141626

Requirement b.

Cost formula = $6899.71 +2.26 per Unit

Cost

No. of cavities

A

High Level

$      52,462.00

20160

B

Low Level

$      25,884.00

8400

C=A-B

Difference

$      26,578.00

11760

A

Cost difference

$      26,578.00

B

No. of cavities difference

11760

C=A/B

Variable cost per unit

$                2.26

High Level

Low Level

A

No. of cavities

20160

8400

B

Variable cost per unit

$                2.26

$             2.26

C=AxB

Total Variable cost

$      45,562.29

$ 18,984.29

D

Total cost

$      52,462.00

$ 25,884.00

E=D-C

Total fixed cost

$        6,899.71

$     6,899.71

Requirement c.

Production level in units

12676

Variable cost

$     36,760.40

Fixed cost

$     30,700.00

Mixed cost

$     35,547.90

Total cost

$ 1,03,008.30